OpenStudy (anonymous):

x−y=4.5 5x+2y=5 how to solve...? and what are these types of questions called

OpenStudy (anonymous):

x = (4.5+y) Substitute this in the place of x for the second equation. You will get the answer.

OpenStudy (anonymous):

:< lemme try 5(4.5+y)+2y=5 22.5+5y+2y=5 7y=-17.5 y=-2.5 x=2 ?

OpenStudy (anonymous):

Yes it's right. 2-(-2.5) = 4.5

OpenStudy (anonymous):

so so... how to solve this? :L \[\log_{a}(x/y)=4.5\] and \[\log_{a}x^{5}y^{2}=5\], find the value of \[\log_{a}\sqrt{xy}\]